1. XYZ Company’s stock is selling for $22.50 per share. It has just announced that it will pay a $1.35 dividend later this year. Dividends are expected to grow 4.5% per year in the future. What rate of return are investors expecting to earn with an investment in XYZ?

2. Company J expects to have free cash flow (FCF) this year of $1,000,000. FCF is expected to grow at a constant rate of 3%. If company J has an 8% discount rate, debt of $15,000,000 and 400,000 shares; what is the value per share?
Answer #1

Part 1:

Rate of return expected is 10.50% as per Dividend Discount model as follows:

в с Formula D Cell reference Values 1 2 3 Groiwth rate of dividend 8 Given 4.50% 4 Next Dividend D1 Given $ 1.35 5 Current Pr

Part 2:

Value per share= Equity value/ Number of shares.

Equity value= Firm value- Debt

Firm Value = Free Cash Flow (FCF) next year/(WACC-g)

Where WACC= Weighted cost of capital or the required discount rate and g= constant growth rate of cash flows.

Given, Free Cash Flow this year (FCF0)=$1,000,000 and growth rate=3%

Hence FCF next year= $1,000,000*(1+3%)= $ 1,030,000

Discount rate =8%

Therefore, Firm Value= $ 1,030,000/(8%-3%) = $ 20,600,000

Given that Debt of the company= $15,000,000 and Number of shares=400,000

Therefore, Value per share= ($ 20,600,000-$15,000,000)/400,000 = $14.00